# From Surveys to Narratives: Rethinking Cultural Value Adaptation in LLMs

### v1

@@ Abstract
Adapting cultural values in Large Language Models (LLMs) presents significant challenges, particularly due to biases and limited training data. Prior work primarily aligns LLMs with different cultural values using World Values Survey (WVS) data. However, it remains unclear whether this approach effectively captures cultural nuances or produces distinct cultural representations for various downstream tasks. In this paper, we systematically investigate WVS-based training for cultural value adaptation and find that relying solely on survey data can homogenize cultural norms and interfere with factual knowledge. To investigate these issues, we augment WVS with encyclopedic and scenario-based cultural narratives from Wikipedia and NormAd. While these narratives may have variable effects on downstream tasks, they consistently improve cultural distinctiveness than survey data alone. Our work highlights the inherent complexity of aligning cultural values with the goal of guiding task-specific behavior.
@@ 2.2 Cultural Value Adaptation via Fine-tuning
Beyond zero-shot prompts, we explore explicit fine-tuning with culture-specific data, referred to as single-culture adaptation in our paper. Following Li et al. (2024a), we train a separate LoRA adapter Hu et al. (2022) for each cultural context using data from a single or a combination of data sources. Each adapter is specialized to reflect the norms, attitudes, or knowledge of that specific culture. However, data sparsity and overfitting are risks, particularly for cultures with limited samples.
@@ 3.4 Models and Training
In this work, we evaluate three variants of LLMs, including Llama-3.1-8B (base and instruction-tuned, Touvron et al. 2023; Dubey et al. 2024), Gemma-2-9B (instruction-tuned, Rivière et al. 2024), and Qwen-2.5-7B (instruction-tuned, Team 2024). In our experiments, all instruction-tuned models are suffixed with “-IT”. We perform LoRA adaptation Hu et al. (2022) on each model using rank-64 LoRA matrices, a batch size of 32, a learning rate of $2\times 10^{-4}$, and six training epochs. Other details on training are in Appendix B.
@@ 3.5 Evaluation Metrics
In the formula above, we normalize by column (i.e., by the test culture) since each test culture set may have different difficulty and scales. This normalization also helps identify which adapter performs best for a given culture.
In an ideal scenario, the best performing adapted model for a particular culture should be based on its own culture, resulting in a C-Dist score of $1.0$. A lower score suggests interference or homogenization, as illustrated in Figure 2. This metric thus quantifies the extent to which each model preserves distinct cultural representations after adaptation.
@@ 4.3 Factual Knowledge Interference
Llama-3.1-8B-IT WVS 0.76 29.61 42.78
Wiki 0.81 35.39 26.33
NormAd 0.85 38.42 19.63
WVS+Wiki 0.78 31.19 49.02
WVS+NormAd 0.89 40.94 50.43
WVS+Wiki+NormAd 0.76 38.21 52.61
Gemma-2-9B-IT WVS 0.81 39.22 45.31
Wiki 0.83 36.67 8.23
NormAd 0.79 37.10 8.07
WVS+Wiki 0.80 37.25 47.05
WVS+NormAd 0.83 40.01 55.19
WVS+Wiki+NormAd 0.73 37.90 64.94
Qwen2.5-7B-IT WVS 0.92 48.05 68.32
Wiki 0.89 44.21 58.32
NormAd 0.91 48.31 65.57
WVS+Wiki 0.90 46.00 68.22
WVS+NormAd 0.94 47.67 67.51
WVS+Wiki+NormAd 0.86 44.13 67.33
Table 5: Averaged performances on the multilingual multicultural offensiveness classifications (F1-Cult.), MMLU evaluations (F1-MMLU), and C-Dist for various instruction-tuned models and data configurations. Augmenting training with NormAd consistently improves C-Dist , but degrades MMLU performance in Llama-3.1-8B-IT and Gemma-2-9B-IT (likely due to reduced instruction-following ability, see Appendix C.2). The results highlight the complexity of adapting cultural values while maintaining cultural distinctiveness, culture-related task performance, and knowledge retention.
@@ 5 Adaptation with Additional Narratives
While WVS-based training provides a great foundation for cultural value adaptation, our results in Section 4 show that this seldom produces strong diagonal patterns, indicating limited cultural specialization. A critical question is what additional data could enhance cultural value adaptation and preserve cultural distinctiveness?
Humans exhibit gaps in what they “think”, and how they “behave” (Gross and Niman, 1975; Fazio, 1981, inter alia). This suggests that self-reported value data, such as the WVS, may be insufficient for improving tasks that require behavioral changes based on cultural values (also evident in our analysis in §4.1). Hence, we incorporate two additional data sources, Wikipedia and NormAd, hypothesizing that introducing data containing more objective narratives of culture could enhance the model’s performance and understanding of cultural values.
@@
Improved C-Dist with Wikipedia and NormAd. The addition of Wikipedia and NormAd significantly enhances cultural distinctiveness (i.e., C-Dist ). Table 5 shows that integrating these datasets consistently improves C-Dist scores across all three models, indicating more culturally distinct behavior. For instance, Llama-3.1-8B-IT’s C-Dist improves from $0.76$ (WVS-only) to $0.89$ (WVS+NormAd). Figure 3 illustrates this shift, as the heatmaps become more diagonal and show reduced cross-cultural interference. Incorporating additional cultural narratives retains cultural distinctiveness.
@@ 6 Further Analysis
Our empirical results suggest that adding objective cultural descriptions and context-specific examples improves cultural distinctiveness and performance on downstream tasks. In this section, we analyze the data further to understand why.
@@ 7 Related Work
Recent research also explored value prediction with In-Context Learning (ICL)-based adaptation methods Choenni and Shutova (2024); Jiang et al. (2024); Myung et al. (2025). Particularly, Jiang et al. (2024) showed a mild inconsistency when models adapted using individual data from one continent were evaluated using data from another (e.g., training data for other continents generally improves alignment to Oceania people). While related to our work, we focus on the impact at the country level rather than the broader continent level.
@@ 5. Offense towards an other (non-human) entity
[…]
@@ F.2 Test Data Statistics
Arabic (Method-Ar) Middle East Offensive language detection: OffensEval2020(2000) (Zampieri et al., 2020b), OSACT4(1000) (Husain, 2020), Multi-Platform(1000) (Chowdhury et al., 2020), and OSACT5(2541) (Mubarak et al., 2022). Hate detection: OSACT4(1000) (Husain, 2020), Multi-Platform(675) (Chowdhury et al., 2020), OSACT5(2541) (Mubarak et al., 2022), and OSACT5_finegrained(2541) (Mubarak et al., 2022). Spam detection: ASHT(1000) (Kaddoura and Henno, 2024). Vulgar detection: Multi-Platform(675) (Chowdhury et al., 2020) 14,973
Bangli (Method-Bn) Bangladesh Offensive language detection: TRAC2020 Task1(1000) (Bhattacharya et al., 2020), TRAC2020 Task2(1000) (Bhattacharya et al., 2020), BAD(1000) (Sharif and Hoque, 2022). Hate detection: Hate Speech(1000) (Romim et al., 2021). Threat detection: BACD(1000) (aimansnigdha, 2018). Bias detection: BACD(1000) (aimansnigdha, 2018). 6,000
Chinese (Method-Zh) China Spam detection: CCS(1000) (Jiang et al., 2019). Bias detection: CDial-Bias(1000) (Zhou et al., 2022). Stance detection: CValues(1712) (Xu et al., 2023). 3,712
English (Method-En) United States Offensive language detection: SOLID(1000) (Rosenthal et al., 2020). Hate detection: MLMA(1000) (Ousidhoum et al., 2019) and HOF(1000) (Davidson et al., 2017). Threat detection: CValuesJMT(1000) (Kaggle, 2019). Toxicity detection: MLMA(1000) (Ousidhoum et al., 2019) and JMT(1000) (Kaggle, 2019). 6,000
German (Method-De) Germany and parts of Europe Offensive language detection: GermEval2018(3531) (Wiegand et al., 2018). Hate detection: IWG_1(469) (Ross et al., 2016), IWG_2(469) (Ross et al., 2016), HASOC2020(850) (HASOC, 2020), and multilingual-hatecheck(1000) (Röttger et al., 2022). 6,319
Korean (Method-Ko) South Korea Hate detection: K-MHaS(1000) (Lee et al., 2022), hateSpeech(1000) (Moon et al., 2020), and HateSpeech2(1000) (daanVeer, 2020). Abusive detection: AbuseEval(1000) (Caselli et al., 2020), CADD(1000) (Song et al., 2021), and Waseem(1000) (Waseem and Hovy, 2016). 5,000
Portuguese (Method-Pt) Brazil and parts of Latin America Offensive language detection: OffComBR(1250) (de Pelle and Moreira, 2017), and HateBR(1000) (Vargas et al., 2022). Bias detection: ToLD-Br-homophobia(1000) (Leite et al., 2020), and ToLD-Br-misogyny(1000) (Leite et al., 2020). Abusive detection: ToLD-Br-insult(1000) (Leite et al., 2020). 16,250
Spanish (Method-Es) Argentina, Mexico, and parts of Latin America Offensive language detection: AMI(1000) (Fersini et al., 2018), MEX-A3T(1000) (Álvarez-Carmona et al., 2018), and OffendES(1000) (Plaza-del Arco et al., 2021). Hate detection: HatEval 2019(1000) (Basile et al., 2019), and HaterNet(1000) (Pereira-Kohatsu et al., 2019). Bias detection: DETOXIS_stereotype(1000) (de Paula and Schlicht, 2021), and DETOXIS_improper(1000) (de Paula and Schlicht, 2021). Abusive detection: DETOXIS_abusive(1000) (de Paula and Schlicht, 2021), DETOXIS_mockery(1000) (de Paula and Schlicht, 2021). Aggressiveness detection: DETOXIS_aggressiveness(1000) (de Paula and Schlicht, 2021). Stance detection: DETOXIS_stance(1000) (de Paula and Schlicht, 2021). 11,000
Turkish (Method-Tr) Turkey Offensive language detection: SemEval-2020(3528) (Zampieri et al., 2020b), offenseCorpus(1000) (Çöltekin, 2020), offenseKaggle(1000) (Kaggle, 2021), and offenseKaggle_2(1000) (Kaggle, 2022). Abusive detection: ATC(1000) (Karayiğit et al., 2021). Spam detection: Turkish Spam(825) (mis, 2019). Fine-grained offensive detection: offenseCorpus(1000) (Çöltekin, 2020). 10,353

### v2

@@ Abstract
Adapting cultural values in Large Language Models (LLMs) presents significant challenges, particularly due to biases and limited training data. Prior work primarily aligns LLMs with different cultural values using World Values Survey (WVS) data. However, it remains unclear whether this approach effectively captures cultural nuances or produces distinct cultural representations for various downstream tasks. In this paper, we systematically investigate WVS-based training for cultural value adaptation and find that relying solely on survey data can homogenize cultural norms and interfere with factual knowledge. To investigate these issues, we augment WVS with encyclopedic and scenario-based cultural narratives from Wikipedia and NormAd. While these narratives may have variable effects on downstream tasks, they consistently improve cultural distinctiveness than survey data alone. Our work highlights the inherent complexity of aligning cultural values with the goal of guiding task-specific behavior. We release our code at https://github.com/faridlazuarda/from-surveys-to-narratives.
@@ 2.2 Cultural Value Adaptation via Fine-tuning
Beyond zero-shot prompts, we explore explicit fine-tuning with culture-specific data, referred to as single-culture adaptation in our paper. Following Li et al. (2024a), we train a separate LoRA adapter Hu et al. (2022) for each cultural context using data from a single or a combination of data sources. Each adapter is specialized to reflect the norms, attitudes, or knowledge of specific culture. However, data sparsity and overfitting are risks, particularly for cultures with limited samples.
@@ 3.1 Linguistic-Cultural Settings
Terminology.
In this paper, we use culture to denote a linguistic–regional grouping (e.g., “Arab” represented by Iraq and Jordan; “Spanish” by Argentina and Mexico). Country refers to specific national datasets used to instantiate a culture. Language denotes the linguistic form of inputs/outputs (e.g., ara, spa). While related, these are not interchangeable; our experiments condition adapters on culture, evaluate on country-specific test sets, and control for language.
@@ 3.3 Evaluation Dataset
Dataset Balance.
Our training and evaluation sets are imbalanced across cultures and languages (see Appendix F; Tables 18, 19, and 20). For example, NormAd token counts range from 102,705 (Arabic) to 6,784 (Korean), and Wikipedia tokens from 15,632 (English) to 3,662 (Spanish). To mitigate confounding, we (i) report per-culture scores alongside macro-averages, (ii) use the C-Dist normalization that compares each adapter against the best performer for each test culture (column-wise), and (iii) repeat all runs with three seeds and report mean $\pm$ sd. We avoid aggressive rebalancing to preserve culture-specific signal, so some absolute gains may still correlate with data availability; we therefore interpret improvements as gains under realistic resource disparity and revisit residual skew in the Limitations.
@@ 3.4 Models and Training
In this work, we evaluate three variants of LLMs, including Llama-3.1-8B (base and instruction-tuned, Touvron et al. 2023; Dubey et al. 2024), Gemma-2-9B (instruction-tuned, Rivière et al. 2024), and Qwen-2.5-7B (instruction-tuned, Team 2024). In our experiments, all instruction-tuned models are suffixed with “-IT”. We perform LoRA adaptation Hu et al. (2022) on each model using rank-64 LoRA matrices, a batch size of 32, a learning rate of $2\times 10^{-4}$, and six training epochs. All adapters were trained on a single 80GB A100 GPU with bf16 and gradient checkpointing. Other details on training are in Appendix B.
We evaluate three open-weight families, each with instruction-tuned variants, to ensure results generalize across distinct pretraining corpora while remaining reproducible Dubey et al. (2024); Rivière et al. (2024); Team (2024). Restricting capacity to 7 to 9B balances multilingual coverage and compute so that PEFT fits on a single high-memory GPU. We adopt per-culture LoRA adapters instead of full fine-tuning to reduce catastrophic forgetting (tracked with MMLU; §4.3) and enable plug-and-play cultural control. Adapters target q_proj/v_proj with rank 64, dropout 0.1, batch size 32, learning rate $2{\times}10^{-4}$, and six epochs, a setting that preserved knowledge while maintaining cultural distinctiveness.
@@ 3.5 Evaluation Metrics
In the formula above, we normalize by column (i.e., by the test culture) since each test culture set may have different difficulty and scales. This normalization also helps identify which adapter performs best for a given culture. In an ideal scenario, the best performing adapted model for a particular culture should be based on its own culture, resulting in a C-Dist score of $1.0$. A lower score suggests interference or homogenization, as illustrated in Figure 2. This metric thus quantifies the extent to which each model preserves distinct cultural representations after adaptation.
@@ 4.3 Factual Knowledge Interference
Llama-3.1-8B-IT WVS $0.76\pm 0.012$ $29.61\pm 1.91$ $42.78\pm 0.84$
Wiki $0.81\pm 0.009$ $35.39\pm 2.77$ $26.33\pm 2.05$
NormAd $0.85\pm 0.014$ $38.42\pm 2.92$ $19.63\pm 1.38$
WVS+Wiki $0.78\pm 0.010$ $31.19\pm 1.80$ $49.02\pm 0.88$
WVS+NormAd $\mathbf{0.89\pm 0.011}$ $\mathbf{40.94\pm 0.98}$ $\mathbf{50.43\pm 0.93}$
WVS+Wiki+NormAd $0.76\pm 0.013$ $38.21\pm 0.85$ $52.61\pm 0.97$
Gemma-2-9B-IT WVS $0.81\pm 0.011$ $39.22\pm 1.66$ $45.31\pm 0.78$
Wiki $0.83\pm 0.010$ $36.67\pm 2.61$ $8.23\pm 1.92$
NormAd $0.79\pm 0.013$ $37.10\pm 2.64$ $8.07\pm 1.96$
WVS+Wiki $0.80\pm 0.012$ $37.25\pm 1.60$ $47.05\pm 0.82$
WVS+NormAd $\mathbf{0.83\pm 0.012}$ $\mathbf{40.01\pm 0.69}$ $\mathbf{55.19\pm 0.86}$
WVS+Wiki+NormAd $0.73\pm 0.018$ $37.90\pm 0.83$ $64.94\pm 1.01$
Qwen2.5-7B-IT WVS $0.92\pm 0.006$ $48.05\pm 1.41$ $68.32\pm 0.56$
Wiki $0.89\pm 0.008$ $44.21\pm 2.38$ $58.32\pm 1.62$
NormAd $0.91\pm 0.007$ $48.31\pm 2.42$ $65.57\pm 0.58$
WVS+Wiki $0.90\pm 0.007$ $46.00\pm 1.39$ $68.22\pm 1.57$
WVS+NormAd $\mathbf{0.94\pm 0.006}$ $47.67\pm 0.40$ $67.51\pm 0.57$
WVS+Wiki+NormAd $0.86\pm 0.010$ $44.13\pm 0.37$ $67.33\pm 0.98$
Table 5: Mean $\pm$ standard deviation across 3 runs with different seeds. NormAd drives the largest gains in cultural distinctiveness (C-Dist ) and cultural-task F1, especially when paired with WVS (Llama/Gemma). Wikipedia alone yields smaller C-DIST gains but, when combined with WVS, stabilizes or improves MMLU (knowledge retention). Variability (sd) is modest for C-DIST and F1, and slightly higher for narrative-only MMLU rows, consistent with greater instability when training without WVS grounding.
@@ 5 Adaptation with Additional Narratives
While WVS-based training provides a useful starting point for cultural value adaptation, our results in Section 4 show that it rarely produces strong diagonal patterns, suggesting limited cultural specialization. This raises a key question: what types of additional data can strengthen adaptation while preserving cultural distinctiveness?
Decades of work in social psychology highlight a long standing gap between what people report in surveys and how they behave in practice (Gross and Niman, 1975; Fazio, 1981, inter alia). This raises concern on the sufficiency of self-reported values such as those in WVS for improving tasks that depend on culturally grounded behavior (§4.1). To bridge this gap, we add WVS with two complementary narrative sources: Wikipedia and NormAd. Unlike survey responses, these provide more objective, context-rich source of cultural norms, institutions, and everyday practices. We hypothesize that incorporating such narratives of culture helps models learn both abstract values and their behavior.
@@
5.1 Complementary roles of Wikipedia and NormAd
Beyond surveys, our additional sources contribute different kinds of supervision. Wikipedia introduces largely declarative, encyclopedic context about institutions, historical trajectories, and everyday practices. Because this material is descriptive and source-based, it tends to regularize the model toward factual consistency and mitigate catastrophic forgetting during cultural finetuning. By contrast, NormAd supplies procedural, scenario-based norms-roles, obligations, and sanctions expressed in concrete situations, so the training signal is closer to how cultural rules are applied in behavior-sensitive settings.
Empirically (Table 5), combining NormAd on top of WVS (WVS+NormAd) gives the largest gains in cultural distinctiveness and culture-task F1. For example, on Llama-3.1-8B-IT we see C-Dist $0.76{\to}0.89$ and F1-Cult. $29.61{\to}40.94$, along with a clearer diagonal in Figure 3. Adding Wikipedia to WVS (WVS+Wiki) usually keeps or improves MMLU while maintaining cultural separation, and using all three together often gives the best MMLU for Llama/Gemma.
We observe that this happens due to NormAd contains concrete, situation-based rules, so the model learns how to apply norms in decisions. Wikipedia provides broad factual background, which helps keep general knowledge stable and offsets drift from training only on surveys (WVS). As a result, narrative-only fine-tuning can make knowledge less stable (Appendix C.2), and WVS-only adapters tend to blur across cultures (§4.2). Mixing WVS+NormAd restores cultural separation, and adding Wikipedia yields the most reliable balance overall.
@@ 5.1 Complementary roles of Wikipedia and NormAd
Improved C-Dist with Wikipedia and NormAd. The addition of Wikipedia and NormAd often enhances cultural distinctiveness, with improvements statistically significant in most cases (see Table 5). Table 5 shows that integrating these datasets consistently improves C-Dist scores across all three models, indicating more culturally distinct behavior. For instance, Llama-3.1-8B-IT’s C-Dist improves from $0.76$ (WVS-only) to $0.89$ (WVS+NormAd). Figure 3 illustrates this shift, as the heatmaps become more diagonal and show reduced cross-cultural interference. Incorporating additional cultural narratives retains cultural distinctiveness.
@@ 6 Further Analysis
Our empirical results suggest that adding objective cultural descriptions and context-specific examples improves cultural distinctiveness and performance on downstream tasks. We also found that pairing WVS with scenario-based narratives (WVS+NormAd; optionally +Wiki) is a strong default when aiming for culturally distinct behavior without sacrificing factual knowledge. In this section, we analyze the data further to understand why.
@@ 7 Related Work
Recent research also explored value prediction with In-Context Learning (ICL)-based adaptation methods Choenni and Shutova (2024); Jiang et al. (2024); Myung et al. (2025). Particularly, Jiang et al. (2024) showed a mild inconsistency when models adapted using individual data from one continent were evaluated using data from another (e.g., training data for other continents generally improves alignment to Oceania people). Similarly, Beck et al. (2024) introduced sensitivity tests to probe how models respond to culturally charged questions, revealing both strengths and blind spots. These evaluation-oriented studies complement adaptation methods by highlighting where misalignments occur and by motivating the development of more nuanced cultural adaptation strategies. While related to our work, we focus on the impact at the country level rather than the broader continent level.
@@ 8 Conclusion
Acknowledgment
This work has been supported by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This research work has also been funded by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We thank Yongxin Huang, German Ortiz for their feedback on a draft of this paper.
@@ 5. Offense towards an other (non-human) entity
(Note: {entity} is derived from the task name, e.g., bias_on_gender_detect $\rightarrow$ gender bias, etc.)
@@ F.2 Test Data Statistics
Arabic
(Method-Ar)
Middle East
Offensive language detection: OffensEval2020(2000) (Zampieri et al., 2020b),
OSACT4(1000) (Husain, 2020),
Multi-Platform(1000) (Chowdhury et al., 2020),
and OSACT5(2541) (Mubarak et al., 2022).
Hate detection: OSACT4(1000) (Husain, 2020),
Multi-Platform(675) (Chowdhury et al., 2020),
OSACT5(2541) (Mubarak et al., 2022),
and OSACT5_finegrained(2541) (Mubarak et al., 2022).
Spam detection: ASHT(1000) (Kaddoura and Henno, 2024).
Vulgar detection: Multi-Platform(675) (Chowdhury et al., 2020)
14,973
Bangli
(Method-Bn)
Bangladesh
Offensive language detection: TRAC2020 Task1(1000) (Bhattacharya et al., 2020),
TRAC2020 Task2(1000) (Bhattacharya et al., 2020),
BAD(1000) (Sharif and Hoque, 2022).
Hate detection: Hate Speech(1000) (Romim et al., 2021).
Threat detection: BACD(1000) (aimansnigdha, 2018).
Bias detection: BACD(1000) (aimansnigdha, 2018).
6,000
Chinese
(Method-Zh)
China
Spam detection: CCS(1000) (Jiang et al., 2019).
Bias detection: CDial-Bias(1000) (Zhou et al., 2022).
Stance detection: CValues(1712) (Xu et al., 2023).
3,712
English
(Method-En)
United States
Offensive language detection: SOLID(1000) (Rosenthal et al., 2020).
Hate detection: MLMA(1000) (Ousidhoum et al., 2019)
and HOF(1000) (Davidson et al., 2017).
Threat detection: CValuesJMT(1000) (Kaggle, 2019).
Toxicity detection: MLMA(1000) (Ousidhoum et al., 2019)
and JMT(1000) (Kaggle, 2019).
6,000
German
(Method-De)
Germany and
parts of Europe
Offensive language detection: GermEval2018(3531) (Wiegand et al., 2018).
Hate detection: IWG_1(469) (Ross et al., 2016),
IWG_2(469) (Ross et al., 2016), HASOC2020(850) (HASOC, 2020),
and multilingual-hatecheck(1000) (Röttger et al., 2022).
6,319
Korean
(Method-Ko)
South Korea
Hate detection: K-MHaS(1000) (Lee et al., 2022),
hateSpeech(1000) (Moon et al., 2020),
and HateSpeech2(1000) (daanVeer, 2020).
Abusive detection: AbuseEval(1000) (Caselli et al., 2020),
CADD(1000) (Song et al., 2021),
and Waseem(1000) (Waseem and Hovy, 2016).
5,000
Portuguese
(Method-Pt)
Brazil and
parts of
Latin America
Offensive language detection: OffComBR(1250) (de Pelle and Moreira, 2017),
and HateBR(1000) (Vargas et al., 2022).
Bias detection: ToLD-Br-homophobia(1000) (Leite et al., 2020),
and ToLD-Br-misogyny(1000) (Leite et al., 2020).
Abusive detection: ToLD-Br-insult(1000) (Leite et al., 2020).
16,250
Spanish
(Method-Es)
Argentina,
Mexico,
and parts of
Latin America
Offensive language detection: AMI(1000) (Fersini et al., 2018),
MEX-A3T(1000) (Álvarez-Carmona et al., 2018),
and OffendES(1000) (Plaza-del Arco et al., 2021).
Hate detection: HatEval 2019(1000) (Basile et al., 2019),
and HaterNet(1000) (Pereira-Kohatsu et al., 2019).
Bias detection: DETOXIS_stereotype(1000) (de Paula and Schlicht, 2021),
and DETOXIS_improper(1000) (de Paula and Schlicht, 2021).
Abusive detection: DETOXIS_abusive(1000) (de Paula and Schlicht, 2021),
DETOXIS_mockery(1000) (de Paula and Schlicht, 2021).
Aggressiveness detection: DETOXIS_aggressiveness(1000) (de Paula and Schlicht, 2021).
Stance detection: DETOXIS_stance(1000) (de Paula and Schlicht, 2021).
11,000
Turkish
(Method-Tr)
Turkey
Offensive language detection: SemEval-2020(3528) (Zampieri et al., 2020b),
offenseCorpus(1000) (Çöltekin, 2020),
offenseKaggle(1000) (Kaggle, 2021),
and offenseKaggle_2(1000) (Kaggle, 2022).
Abusive detection: ATC(1000) (Karayiğit et al., 2021).
Spam detection: Turkish Spam(825) (mis, 2019).
Fine-grained offensive detection: offenseCorpus(1000) (Çöltekin, 2020).
10,353